OpenStudy (jojo12g):

i think c Subtract. 3 – (–18) A. –21 B. –15 C. 15 D. 21

OpenStudy (alexandervonhumboldt2):

wrong

OpenStudy (alexandervonhumboldt2):

look: -(-)=+

OpenStudy (anonymous):

d

OpenStudy (alexandervonhumboldt2):

thus 3-(-18)=3+18=???

OpenStudy (alexandervonhumboldt2):

please do not give out direct answers. it is against the rules. explain your answer.

OpenStudy (anonymous):

d

OpenStudy (alexandervonhumboldt2):

please do not give out direct answers. it is against the rules. explain your answer.

OpenStudy (anonymous):

http://openstudy.com/images/medal.png

OpenStudy (jojo12g):

is d correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

medal and bestreonse

OpenStudy (jojo12g):

thanks guys both of u

OpenStudy (alexandervonhumboldt2):

please do not ask for medals it is against the rules @jayshon54rogers

OpenStudy (alexandervonhumboldt2):

you are welcome @jojo12g

OpenStudy (ddos_dragon):

Sorry I could not get here sooner. :) Make sure and medal whoever you thought answered your question best and close the question. Have a nice day.